Ayana has $2.35 in nickels and dimes. If she has 33 coins in all, find the number of nickels and dimes

Answer by checkley79(3341)

N+D=33
N=33-D
.05N+.10D=2.35
.05(33-D)+.10D=2.35
1.65-.05D+.10D=2.35
.05D=2.35-1.65
.05D=.70
D=.70/.05
D=14 DIMES.
N=33-14=19 NICKLES.
PROOF:
.05*19+.10*14=2.35
.95+140=2.35
2.35=2.35